OpenStudy (anonymous):

can help me solve integral 1/surt3x dx

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ \sqrt{3x} }dx\]

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

ok , continue

OpenStudy (anonymous):

mimi , u r incorrect

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{3} } \int\limits \frac{ 1 }{ \sqrt{x} }dx => \frac{ 1 }{ \sqrt{3} }\times2\sqrt{x}+C\]

OpenStudy (anonymous):

the ans is 2 surt x/3 plus c

OpenStudy (anonymous):

\[\frac{ 2\sqrt{x} }{ \sqrt{3} }+C\]

OpenStudy (anonymous):

ya batman correct

OpenStudy (anonymous):

batman teach me plz

OpenStudy (anonymous):

Factored out the constant \[\frac{ 1 }{ \sqrt{3} }\]

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

is there anyway I can do without splitting the fracion ?

OpenStudy (anonymous):

And the derivative of \[\frac{ 1 }{ \sqrt{x} } \ is\ 2\sqrt{x}\]

OpenStudy (anonymous):

u ahven answer my question

OpenStudy (anonymous):

*haven

OpenStudy (anonymous):

It'll make your life harder

OpenStudy (anonymous):

You can try it out, but I have no interest.

OpenStudy (anonymous):

lol not helping ...

OpenStudy (anonymous):

I don't know what you want?

OpenStudy (anonymous):

u try solving it without splitting

OpenStudy (anonymous):

I didn't split anything, I just factored out the constant.

OpenStudy (rational):

try \(x = 3 \sin^2 \theta\)

OpenStudy (anonymous):

try do without factorise

OpenStudy (rational):

^if you want to make the simple problem challenging :P

OpenStudy (anonymous):

I want see two method ( my exam will ask )

OpenStudy (anonymous):

Why would I want to make it harder on myself.

OpenStudy (rational):

Ohk, there are plenty of ways to do this actually... let me think of next best method

OpenStudy (anonymous):

I guess you can try what rational mentioned, but I doubt it'll get you anywhere.

OpenStudy (anonymous):

I got wrong for this in exam and Im pissed off , I did without factorizing , I just move the surt 3x to the top and start integrating and put differentiate inside at the bottom ......

OpenStudy (anonymous):